# Generalized Multi-Objective Reinforcement Learning with Envelope Updates in URLLC-enabled Vehicular Networks

### v1

@@ Abstract
We develop a novel multi-objective reinforcement learning (MORL) framework to jointly optimize wireless network selection and autonomous driving policies in a multi-band vehicular network operating on conventional sub-6GHz spectrum and Terahertz frequencies. The proposed framework is designed to (i) maximize the traffic flow and minimize collisions by controlling the vehicle’s motion dynamics (i.e., speed and acceleration), and (ii) enhance the ultra-reliable low-latency communication (URLLC) while minimizing handoffs (HOs). We cast this problem as a multi-objective Markov Decision Process (MOMDP) and develop solutions for both predefined and unknown preferences of the conflicting objectives. Specifically, deep-Q-network and double deep-Q-network-based solutions are developed first that consider scalarizing the transportation and telecommunication rewards using predefined preferences. We then develop a novel envelope MORL solution which develop policies that address multiple objectives with unknown preferences to the agent. While this approach reduces reliance on scalar rewards, policy effectiveness varying with different preferences is a challenge. To address this, we apply a generalized version of the Bellman equation and optimize the convex envelope of multi-objective Q values to learn a unified parametric representation capable of generating optimal policies across all possible preference configurations. Following an initial learning phase, our agent can execute optimal policies under any specified preference or infer preferences from minimal data samples. Numerical results validate the efficacy of the envelope-based MORL solution and demonstrate interesting insights related to the inter-dependency of vehicle motion dynamics, HOs, and the communication data rate. The proposed policies enable autonomous vehicles to adopt safe driving behaviors with improved connectivity.
@@ I Introduction
On the other hand, the use of deep reinforcement learning (DRL) is becoming critical for online decision making in highly random mobility-oriented wireless environments. In the context of V2I communications, a plethora of research works focused on improving network quality of service (QoS) (e.g., including transmission delay, link throughput, etc) via DRL-based resource allocation [2, 3, 4]. These research works were focused on considering sub-channel and power allocation to enhance V2I communications. Particularly, the authors in [2, 3] adopted deep Q-network (DQN) and multi-agent DQN to enhance the total throughput of V2I connections and the payload delivery rate of vehicle-to-vehicle (V2V) connections simultaneously. Xu et al. [4] derived the contribution-based dual-clip proximal policy to optimize V2I and V2V links separately. However, their system model only contains a single BS where handovers (HOs) are not considered.
@@
Along another note, most of the existing research works in the transportation are focused on collision-avoidance [10, 11], safe driving [12], and efficient fuel consumption [12, 13, 14]. For instance, in [10], the authors applied a RL framework for faster travel and the reward is proportional to the AV’s velocity along with a penalty for vehicle collision. The action space included acceleration, deceleration, lane changes (LC), and maintain speed, whereas the state space was based on AVs’ locations and their respective velocities. In [12], the authors applied DDQN to enhance AVs’ driving safety and fuel consumption. The state space included AVs’ locations, fuel consumption, and velocities, whereas the actions included speeds and LC of AVs. In [13], the authors applied the Intelligent Driver Model (IDM) and Minimizing Overall Braking Induced by Lane Change (MOBIL) to control steering and lane change. The proposed reward design encourages long traffic, high speed and discourages unnecessary LC. Authors in [14] introduced multi-objective actor-critic to improve the tradeoff between energy consumption and travel efficiency of AVs. The authors derived MO actor-critic to optimize two objectives.
To date, none of the existing research works [2, 4, 3, 8, 9, 5, 15, 6, 7, 10, 12, 13, 14] have considered the inter-dependency of the AV motion dynamics to wireless data rates. To improve the communication data rate and minimize road collisions of connected vehicles jointly, it is critical to optimize the AVs’ network selection and driving policies simultaneously. In the sequel, our contributions can be summarized as follows:
@@
  • The considered problem is formulated as a multi-objective Markov decision process (MOMDP) that has two-dimensional action space and rewards consist of telecommunication and autonomous driving utilities. We then propose single policy MORL solutions with predefined preferences thus converting the MOOP into a single-objective and apply DQN and double DQN solutions. The resulting optimal policy depends on the relative preferences of the objectives.
  • Learning optimized policies across multiple preferences remains challenging. To address this, we then develop a novel envelope MORL solution to effectively navigate the entire spectrum of preferences within a given domain. This approach empowers the trained model to generate the best possible policy tailored to any user-defined preference. Our algorithm hinges on two fundamental insights: firstly, we demonstrate that the optimality operator governing a generalized Bellman equation with preferences exhibits valid contraction properties. Secondly, by optimizing for the convex envelope of multi-objective Q-values, we ensure an efficient alignment between preferences and the resultant optimal policies. Leveraging hindsight experience replay, we recycle transitions to facilitate learning across various sampled preferences, while employing homotopy optimization to maintain manageable learning processes.
@@
II System Model and Problem Formulation
A multi-band downlink network comprising $n_{R}$ RBSs and $n_{T}$ TBSs is considered. A multi-vehicle network is also considered, with a multi-lane road comprising $N_{L}$ lanes. $M$ AVs receive information from the BSs (deployed alongside the road) through V2I communications (as depicted in Figure 1). Each AV is permitted to associate with only one BS at a time, regardless of whether the BS is an RBS or TBS. The on-board units (OBUs) on the AVs receive real-time information from the VNet, including the velocity, acceleration, and lane position of surrounding vehicles. Each RBS and TBS has a bandwidth available to it, represented by $W_{R}$ and $W_{T}$, respectively. Each RBS and TBS is capable of supporting a maximum number of AVs, denoted by $Q_{R}$ and $Q_{T}$, respectively. All AVs are equipped with a single antenna.
@@ II-A Downlink V2I Data Transmission Model
where $P_{R}^{\mathrm{tx}},G_{R}^{\mathrm{tx}},G_{R}^{\mathrm{rx}},c,f_{R},r_{ij},$ and $\alpha$ denote the transmit power of the RBSs, antenna transmitting gain, antenna receiving gain, speed of light, RF carrier frequency (in GHz), distance between the $j$-th AV and the $i$-th RBS, and path-loss exponent, respectively. In addition, $H_{i}$ is the exponentially distributed channel fading power observed at the $j$-th AV from the $i$-th RBS, $\sigma^{2}$ is the power of thermal noise at the receiver, $I_{R_{j}}$ is the cumulative interference at $j$-th AV from the interfering RBSs, i.e., where $P_{R}^{\text{tx}}$, $G_{R}^{\text{tx}}$, $G_{R}^{\text{rx}}$, $c$, $f_{R}$, $r_{ij}$, and $\alpha$ represent the transmit power of the RBSs, the gain of the transmitting antenna, the gain of the receiving antenna, the speed of light, the RF carrier frequency in GHz, the distance between the $j$-th AV and the $i$-th RBS, and the path-loss exponent, respectively. Furthermore, $H_{i}$ denotes the exponentially distributed channel fading power observed by the $j$-th AV from the $i$-th RBS, $\sigma^{2}$ is the power of the thermal noise at the receiver, and $I_{R_{j}}$ is the cumulative interference experienced by the $j$-th AV from other interfering RBSs. $I_{R_{j}}=\sum_{k\neq i}P_{R}^{\mathrm{tx}}\gamma_{R}r_{kj}^{-\alpha}H_{k}$ where $r_{kj}$ is the distance between the $k$-th interfering RBS and the $j$-th AV, $H_{k}$ is the power of fading from the $k$-th interfering RBS to the $j$-th AV, and $\gamma_{R}=G_{R}^{\mathrm{tx}}\>G_{R}^{\mathrm{rx}}\left({c}/{4\pi f_{R}}\right)^{2}$. In the context of a Terahertz (THz) network, where molecular absorption significantly impacts signal propagation, the significance of line-of-sight (LOS) transmissions over non-line-of-sight (NLOS) transmissions is dominant. Consequently, the SINR for a given $j$-th AV can be modeled as follows [17]:
@@
where $W_{j},L_{B},\epsilon_{c},f_{Q}^{-1}(\cdot),V$ are the transmission bandwidth of BS $j$, blocklength, decoding error probability, the inverse $Q$ function, and the channel dispersion, respectively. $V$ can be calculated by $1-\frac{1}{(1+\mathrm{SINR}_{ij})^{2}}$. Given that $D_{t}$ time to transmit $L_{B}$ symbols, the time and frequency resources can be computed by $D_{t}W=L_{B}$. As block length $L_{B}$ approaches infinity, the achieve rate in (5) reaches Shannon’s capacity.
@@ II-B Transportation Model
II-B1 Kinematics Model for all AVs
Following [20, 16], we update all AVs’ real-time physical location based on the Kinematics model. Suppose only front wheels can be steered, we have the following relations applied for each AV $j$, $j\in M$:
$\frac{\partial}{\partial t}(x_{j})=v_{j}\cos(\psi_{j}+\beta_{j}),\quad\beta_{j}=\arctan\left({\frac{\tan{\delta_{j}^{\mathrm{fa}}}}{2}}\right)$ (8)
$\frac{\partial}{\partial t}(y_{j})=v_{j}\sin(\psi_{j}+\beta_{j})$ (9)
$\frac{\partial}{\partial t}(v_{j})=a_{j},\quad\frac{\partial}{\partial t}(\psi_{j})=\frac{v_{j}}{l_{j}}\sin{\beta_{j}}$ (10)
where $(x_{j},y_{j})$ is the location of AV $j$. $\psi_{j},a_{j},\beta_{j},l_{j},\delta_{j}^{\mathrm{fa}}$ are AV $j$’s heading, command for acceleration, slip angle at the gravity center, half-length of AV $j$, and front wheel angle, respectively.
II-B2 Acceleration and Lane Change of Target Vehicles
As in [16, 21], each target AV $j$ follow a lane $L_{j}$ through its lateral position $y_{j}$ and heading $\psi_{L_{j}}$, for $i\in M_{1}$. Thus, target AVs follow a cascade controller of lateral position and heading, i.e.,
$\frac{\partial}{\partial t}(\psi_{j})=K_{j}^{\psi}\left[(\psi_{L_{j}}+\arcsin\left(\frac{\tilde{v}_{i,y}}{v_{j}}\right)-\psi_{j}\right]$ (11)
where $\tilde{v}_{i,y}=K_{j}^{y}\left(y_{L_{j}}-y_{j}\right),$ $K_{j}^{\psi}$ and $K_{j}^{y}$ are the control gains. $y_{L_{j}}$ is the lane $L_{j}$ lateral position. The acceleration of target AV $j$, for $j\in M_{1}$, is given as[16, 21]:
$a_{j}=K_{0}^{v}(v_{r}-v_{j})$ (12)
where $v_{r}$ is the desired AV speed and $K_{0}^{v}$ is the control gain.
II-B3 Acceleration and Lane Change of Surrounding Vehicles
Note that surrounding AVs will not get involved in training. Instead, the surrounding AVs select their accelerations/decelerations based on the IDM and MOBIL model, which allows them to track for a specific target speed and follow in a specific lane. For each surrounding vehicle, we compute the acceleration using the Intelligent Driver Model (IDM) [16, 22] and applied the MOBIL model for changing lanes [23, 16]. IDM is based on the idea that surrounding AVs choose their acceleration or deceleration based on their current speed and the distance to the vehicle in front. The model combines a desire to drive at a certain free-flow speed with a reaction to the traffic situation, particularly to avoid collisions. We define that AV $j$ own velocity $v_{j}$, a real gap distance $d_{j}$ to the leading AV, and the speed difference $\Delta v_{j}$ between two AVs. With IDM, the acceleration $a_{j}$ and desired minimum gap $\hat{d_{j}}$ from the vehicle ahead of AV $j$ , $j\in M_{2}$ is modeled as [16, 22]:
$a_{j}=a_{c}-a_{c}\left[\left(\frac{|v_{j}|}{v_{0}}\right)^{\delta_{a}}+\left(\frac{\hat{d_{j}}}{d_{j}}\right)^{2}\right],$ (13)
where
@@ II-B3 Acceleration and Lane Change of Surrounding Vehicles
where $a_{c},b_{c},v_{0}$ are predefined parameters denoting AV’s maximum acceleration, comfortable braking deceleration, and desired velocity, respectively. $\delta_{a}$ denotes acceleration reduction factor, i.e., a higher $\delta_{a}$ reduces acceleration. $d_{0}$ and $T$ denote the minimum distance in stopped traffic and safe time to approach the front vehicle, respectively. The desired gap has a steady-state (equilibrium) term and a dynamic term that implements the intelligent braking strategy.
To maintain a high transportation efficiency, redundant brakes during lane changing for surrounding AVs are discouraged. MOBIL model focuses on minimizing overall braking induced by a lane change in AV lateral behavior. MOBIL model indicates that an AV should change lanes when one of the following conditions is met, (1) AVs can accelerate more after changing lane $\tilde{a}_{j}\geq-b_{\textbf{safe}},$ where $b_{\textbf{safe}}$ shows the maximum braking imposed on AV when the AV cuts into the adjacent lane, (2) vehicle imposes unsafe and incentives braking on their new following vehicle, i.e,
$\tilde{a_{j}}-a_{j}+p\left(\tilde{a_{o}}-a_{o}+\tilde{a_{n}}-a_{n}\right)\geq\Delta a_{\textbf{th}}$ (15)
where $a_{j}$ and $\tilde{a}_{j}$ are the acceleration of AV before and after lane change. The subscript $o$ and $n$ denote the AV’s older follower and the new follower before and after the lane change, respectively. $p$ represents the politeness coefficient. $\Delta a_{\textbf{th}}$ is the acceleration gain to execute the lane change for AV. A decision of lane change adjusts the target lane $L_{j}$ of AV $j$ on its current lane segment. The actual motion planning and steering to track the newly targeted lane are then executed by the lateral controller in (11).

### v2

@@ Abstract
We develop a novel multi-objective reinforcement learning (MORL) framework to jointly optimize wireless network selection and autonomous driving policies in a multi-band vehicular network operating on conventional sub-6GHz spectrum and Terahertz frequencies. The proposed framework is designed to (i) maximize the traffic flow and minimize collisions by controlling the vehicle’s motion dynamics (i.e., speed and acceleration), and (ii) enhance the ultra-reliable low-latency communication (URLLC) while minimizing handoffs (HOs). We cast this problem as a multi-objective Markov Decision Process (MOMDP) and develop solutions for both predefined and unknown preferences of the conflicting objectives. Specifically, deep-Q-network and double deep-Q-network-based solutions are developed first that consider scalarizing the transportation and telecommunication rewards using predefined preferences. We then develop a novel envelope MORL solution which develop policies that address multiple objectives with unknown preferences to the agent. While this approach reduces reliance on scalar rewards, policy effectiveness varying with different preferences is a challenge. To address this, we apply a generalized version of the Bellman equation and optimize the convex envelope of multi-objective Q values to learn a unified parametric representation capable of generating optimal policies across all possible preference configurations. Following an initial learning phase, our agent can execute optimal policies under any specified preference or infer preferences from minimal data samples. Numerical results validate the efficacy of the envelope-based MORL solution and demonstrate interesting insights related to the inter-dependency of vehicle motion dynamics, HOs, and the communication data rate. The proposed policies enable autonomous vehicles (AVs) to adopt safe driving behaviors with improved connectivity.
@@ I Introduction
On the other hand, the use of Deep Reinforcement Learning (DRL) is becoming critical for online decision making in highly random, mobile-oriented wireless environments. In the context of V2I communications, a wealth of research has focused on improving network quality of service (QoS) (e.g., including transmission delay, link throughput, etc.) via DRL-based resource allocation [2, 3, 4]. This research has focused on considering subchannel and power allocation to improve V2I communication. In particular, the authors in [2, 3] adopted deep-Q network (DQN) and multi-agent DQN to simultaneously improve the overall throughput of V2I links and the payload delivery rate of vehicle-to-vehicle (V2V) links. Xu et al. [4] derived the contribution-based dual-clip proximal policy to optimise V2I and V2V connections separately. However, their system model includes only a single BS where handovers (HOs) are not considered.
@@
Along another note, most of the existing research works in the transportation are focused on collision-avoidance [10, 11], safe driving [12], and efficient fuel consumption [12, 13, 14]. For instance, in [10], the authors applied a RL framework for faster travel and the reward is proportional to the AV’s velocity along with a penalty for vehicle collision. The action space included acceleration, deceleration, lane changes (LC), and maintain speed, whereas the state space was based on AVs’ locations and their respective velocities. In [12], the authors applied DDQN to enhance AVs’ driving safety and fuel consumption. The state space included AVs’ locations, fuel consumption, and velocities, whereas the actions included speeds and LC of AVs. In [13], the authors applied the Intelligent Driver Model (IDM) and Minimizing Overall Braking Induced by Lane Change (MOBIL) to control steering and lane change. The proposed reward design encourages long traffic, high speed and discourages unnecessary LC. Authors in [14] introduced multi-objective actor-critic to improve the tradeoff between energy consumption and travel efficiency of AVs. The authors derived MO actor-critic to optimize two objectives. Recently, a few optimization-based research works have started looking into this problem where the AV velocity has been optimized using simplified car-following models [15, 16, 17, 18, 19, 20], but these models are computationally expensive, do not learn from past experiences thus not relevant for faster decision-making, and are not applicable to dynamic traffic scenarios with various control and communication parameters.
Ref. Autonomous Driving Vehicular Communication ML Optimization Method Computational Complexity
Collision
Avoidance
Speed
Management
Driving
Behaviour
Lane
Changes
Mobility Aware
URLLC
Handover
Management
Interference
Aware
Network
Availability
[1] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
RF + THz
Multibands
Q + DQN Low
[12] $\checkmark$ $\checkmark$ $\checkmark$ AGWN DQN + DDQN Low
[2] $\checkmark$
Single Cell
system 2 GHz
DQN Low
[4] $\checkmark$ $\checkmark$
Multi-Platoon
Vehicular
MORL: CD-PPO High
[14] $\checkmark$ $\checkmark$ $\checkmark$ RF MORL: MOAC High
[21] $\checkmark$ $\checkmark$
UAV Assisted
MEC
EMORL High
[22] $\checkmark$ $\checkmark$ $\checkmark$ AGWN
MORL: thresholded
lexicographic
High
This paper $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
RF + THz
Multibands
MORL
MO-DDQN-Envelope
Low
Table I: Comparison between Related Works and This Work
To date, none of the existing research works [2, 4, 3, 8, 9, 5, 23, 6, 7, 10, 12, 13, 14, 15, 16, 17] have considered the inter-dependency of the AV motion dynamics to wireless data rates.
Different from the existing research, our contributions can be explained from two perspectives, i.e., (1) jointly optimizing both autonomous driving policies and telecommunication rewards with and without predefined preferences, and (2) proposing a novel multi-objective MO-DDQN-Envelope reinforcement learning solution that enables optimizing policies across varying preferences (i.e., without predefined preferences) in the multi-objective domain. Table I summarizes the distinction of existing research in terms of computational complexity, methodology, and performance metrics. Our contributions can be summarized as follows:
@@
  • The considered problem is formulated as a multi-objective Markov decision process (MOMDP) that has two-dimensional action space and rewards consist of telecommunication and autonomous driving utilities. We jointly optimizing both autonomous driving policies and telecommunication rewards with and without predefined preferences using multi-objective reinforcement learning (MORL) approaches.
  • We propose a novel multi-objective MO-DDQN-Envelope reinforcement learning solution that enables optimizing policies across varying preferences (i.e., without predefined preferences) in the multi-objective domain. Our framework simultaneously optimizes telecommunication and transportation objectives in dynamic environments, balancing trade-offs like collision avoidance, velocity management, handover optimization, and network availability. Unlike scalarized methods, our MO-DDQN-Envelope approach dynamically adjusts preferences, mitigating biases and errors introduced by fixed weightings.
@@
II System Model and Assumptions
A multi-band downlink network comprising $n_{R}$ Radio Frequency Base Stations (RBSs) and $n_{T}$ Terahertz Frequency Base Stations (TBSs) is considered. A multi-vehicle network is also considered, with a multi-lane road comprising $N_{L}$ lanes. $M$ AVs receive information from the BSs (deployed alongside the road) through V2I communications (as depicted in Figure 1). Each AV is permitted to associate with only one BS at a time, regardless of whether the BS is an RBS or TBS. The on-board units (OBUs) on the AVs receive real-time information from the VNet, including the velocity, acceleration, and lane position of surrounding vehicles. Each RBS and TBS has a bandwidth available to it, represented by $W_{R}$ and $W_{T}$, respectively. Each RBS and TBS is capable of supporting a maximum number of AVs, denoted by $Q_{R}$ and $Q_{T}$, respectively. All AVs are equipped with a single antenna.
@@ II-A Downlink V2I Data Transmission Model
where $P_{R}^{\text{tx}}$, $G_{R}^{\text{tx}}$, $G_{R}^{\text{rx}}$, $c$, $f_{R}$, $r_{ij}$, and $\alpha$ represent the transmit power of the RBSs, the gain of the transmitting antenna, the gain of the receiving antenna, the speed of light, the RF carrier frequency in GHz, the distance between the $j$-th AV and the $i$-th RBS, and the path-loss exponent, respectively. Note that $r_{ij}={(d_{ij}^{2}+h_{ij}^{2})}^{1/2},$ where $d_{ij}$ is the 2D distance between the $j$-th AV and $i$-th BS and $h_{ij}$ is the transmit antenna height. Furthermore, $H_{i}$ denotes the exponentially distributed channel fading power observed by the $j$-th AV from the $i$-th RBS, $\sigma^{2}$ is the power of the thermal noise at the receiver, and $I_{R_{j}}$ is the cumulative interference experienced by the $j$-th AV from other interfering RBSs. $I_{R_{j}}=\sum_{k\neq i}P_{R}^{\mathrm{tx}}\gamma_{R}r_{kj}^{-\alpha}H_{k}$ where $r_{kj}$ is the distance between the $k$-th interfering RBS and the $j$-th AV, $H_{k}$ is the power of fading from the $k$-th interfering RBS to the $j$-th AV, and $\gamma_{R}=G_{R}^{\mathrm{tx}}\>G_{R}^{\mathrm{rx}}\left({c}/{4\pi f_{R}}\right)^{2}$.
In the context of a Terahertz (THz) network, where molecular absorption significantly impacts signal propagation, the significance of line-of-sight (LOS) transmissions over non-line-of-sight (NLOS) transmissions is dominant. Consequently, the SINR for a given $j$-th AV can be modeled as follows [25]:
@@
where $W_{j},L_{B},\epsilon_{c},f_{Q}^{-1}(\cdot),V$ are the transmission bandwidth of BS $j$, blocklength, decoding error probability, the inverse $Q$ function, and the channel dispersion, respectively. $V$ can be calculated by $1-\frac{1}{(1+\mathrm{SINR}_{ij})^{2}}$. Given that $D_{t}$ time to transmit $L_{B}$ symbols, the time and frequency resources can be computed by $D_{t}W=L_{B}$. where $W=W_{R}$ for RBSs and $W=W_{T}$ for TBSs. As the block length $L_{B}$ approaches infinity, the achieve rate in (5) reaches Shannon’s capacity.
@@ II-B Transportation Model
Following [28, 24], we update the real-time physical location of all AVs using the Kinematics model [28]. Assuming only the front wheels can be steered, the dynamics of each AV $j$, $j\in M$, are described as:
$\frac{\partial x_{j}}{\partial t}=v_{j}\cos(\psi_{j}+\beta_{j}),\quad\frac{\partial y_{j}}{\partial t}=v_{j}\sin(\psi_{j}+\beta_{j}),\quad\frac{\partial v_{j}}{\partial t}=a_{j},$ (8)
where $(x_{j},y_{j})$ represents the position of AV $j$, $v_{j}$ its velocity, $a_{j}$ its commanded acceleration, $\psi_{j}$ its heading angle, and $\beta_{j}$ the slip angle at the vehicle’s center of gravity. The slip angle $\beta_{j}$ is given by:
$\beta_{j}=\arctan\left(\frac{\tan\delta_{j}^{\mathrm{fa}}}{2}\right),$ (9)
where $\delta_{j}^{\mathrm{fa}}$ is the steering angle of the front wheels.
The heading dynamics of AV $j$ depend on its control model:
$\frac{\partial\psi_{j}}{\partial t}=\begin{cases}K_{j}^{\psi}\left[\psi_{L_{j}}+\arcsin\left(\frac{\tilde{v}_{j,y}}{v_{j}}\right)-\psi_{j}\right],&\text{if }j\in M_{1},\\ \frac{v_{j}}{l_{j}}\sin\beta_{j},&\text{if }j\in M_{2},\end{cases}$ (10)
where $l_{j}$ is the half-length of AV $j$, $\psi_{L_{j}}$ is the desired heading, and $K_{j}^{\psi}$ and $K_{j}^{y}$ are control gains. The term $\tilde{v}_{j,y}$, representing lateral velocity adjustment, is given as:
$\tilde{v}_{j,y}=K_{j}^{y}(y_{L_{j}}-y_{j}),$
with $y_{L_{j}}$ being the lateral position of the desired lane.
The acceleration $a_{j}$ for AV $j$ is defined as:
$a_{j}=\begin{cases}K_{0}^{v}(v_{r}-v_{j}),&\text{if }j\in M_{1},\\ a_{c}-a_{c}\left[\left(\frac{|v_{j}|}{v_{0}}\right)^{\delta_{a}}+\left(\frac{\hat{d}_{j}}{d_{j}}\right)^{2}\right],&\text{if }j\in M_{2},\end{cases}$ (11)
where $v_{r}$ is the desired speed, $K_{0}^{v}$ is the speed control gain, $a_{c}$ is the maximum acceleration, $v_{0}$ is the desired velocity, $\delta_{a}$ is the acceleration reduction factor, and $d_{j}$ is the distance to the front AV. The desired gap $\hat{d}_{j}$ is computed as:
@@
where $d_{0}$ is the minimum distance in stopped traffic, $T$ is the safe time gap, $\Delta v_{j}$ is the relative velocity to the front vehicle, and $b_{c}$ is the comfortable braking deceleration.
For AVs in $M_{2}$, discrete lane changes are determined by the MOBIL model [29, 30], where an AV decides to change lanes when (1) AVs are safe to cut-in another lane as $\tilde{a}_{j}\geq-b_{\text{safe}}$. (2) there is an incentive benefit if the target AVs and followers satisfies $\tilde{a}_{j}-a_{j}+p\left(\tilde{a}_{o}-a_{o}+\tilde{a}_{n}-a_{n}\right)\geq\Delta a_{\text{th}}$. where $-b_{\text{safe}}$ is the maximum braking imposed on the new following AV, $p$ is the politeness coefficient, $\Delta a_{\text{th}}$ is the minimum acceleration gain, and subscripts $o$ and $n$ denote the old and new followers, respectively.